OpenStudy (anonymous):

x+3 over 5+x over 3 equals 7

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

LCM is 15

OpenStudy (anonymous):

can you show me how the problem is worked

OpenStudy (anonymous):

you got the right answer

OpenStudy (anonymous):

Yifan your answr is right I just have to show the work can you show me

OpenStudy (anonymous):

3(x+3)+5x=105 8x=96 x=12

OpenStudy (anonymous):

thank you so much I might have one more for you what do you do with the seven in that last problem

OpenStudy (anonymous):

?which problem

OpenStudy (anonymous):

the x=12 what happens with the =7

OpenStudy (anonymous):

new problem v=4/3pie r squared h solve for h (this is Geometry)

OpenStudy (anonymous):

7*15=105

OpenStudy (anonymous):

h=3v/(4pi r squared)

OpenStudy (anonymous):

can you show me how to get that one too please

OpenStudy (anonymous):

treat 4/3 pi r squared as a whole

OpenStudy (anonymous):

thank you and I hope I can find you again when I need you!!!